# Adaptive Paradigm Synergy: Can a Cross-Paradigm Objective Enhance Long-Tailed Learning?

### v1

@@ Abstract
Self-supervised learning (SSL) has achieved impressive results across several computer vision tasks, even rivaling supervised methods. However, its performance degrades on real-world datasets with long-tailed distributions due to difficulties in capturing inherent class imbalances. Although supervised long-tailed learning offers significant insights, the absence of labels in SSL prevents direct transfer of these strategies. To bridge this gap, we introduce Adaptive Paradigm Synergy (APS), a cross-paradigm objective that seeks to unify the strengths of both paradigms. Our approach reexamines contrastive learning from a spatial structure perspective, dynamically adjusting the uniformity of latent space structure through adaptive temperature tuning. Furthermore, we draw on a re-weighting strategy from supervised learning to compensate for the shortcomings of temperature adjustment in explicit quantity perception. Extensive experiments on commonly used long-tailed datasets demonstrate that APS improves performance effectively and efficiently. Our findings reveal the potential for deeper integration between supervised and self-supervised learning, paving the way for robust models that handle real-world class imbalance.
@@ 1 Introduction
Figure 1: Comparison between SimCLR[6] baseline trained on balanced CIFAR10[27] and imbalanced CIFAR10-LT[12] and our APS trained on CIFAR10-LT. Three Splits (Head, Mid and Tail) are shown as different colours (blue, red and green) respectively[31]. We also report the average accuracy and range of variation over ten classes.
@@
Self-supervised learning (SSL)[19, 18] that aims to learn a generic robust feature extractor from unlabeled data has been consistently explored in recent year, achieving results comparable to those of supervised pre-training [8, 4, 47]. Contrastive learning, a leading SSL paradigm, seeks to learn discriminative representations by encouraging the similarity of augmented view pairs, and has achieved remarkable success in computer vision (CV), natural language processing (NLP) and graph learning (GL) [48, 16]. Nevertheless, most contrastive learning methods are conducted on curated datasets like ImageNet[13]. In contrast, real-world data distributions are mainly characterized by long-tailed distributions [36], which could cause a significant decline in performance, as illustrated in Fig. 1. On the relatively simple CIFAR10-LT dataset, the baseline model demonstrates a notable performance collapse, with an average drop exceeding 10% compared to training on the balanced version. Furthermore, for certain tail classes, the accuracy even falls below 30%, indicating SSL still suffers from long-tailed problem. Meanwhile, such self-supervised training pattern of only unlabeled data also makes it hard to rectify or even detect this imbalance. The difficulty of detecting or modeling such inherent implicit imbalance is a significant obstacle to migrating contrastive learning to large-scale, real-world datasets.
In terms of long-tailed learning, several aspects have been explored in recent years [2, 52, 38, 20], e.g., balanced loss function, multi-experts ensemble learning and data augmentation. Extensive experiments on multiple publicly available long-tailed datasets demonstrate the effectiveness of these methods in mitigating performance degradation caused by tail class bias. However, most of the existing research continues to focus on the supervised learning paradigm, with limited attention paid to self-supervised long-tailed recognition. Although several works [22, 54, 53] have been proposed to address long-tailed issues in self-supervised settings, they often treat this as a separate domain and rarely draw on the experiments of supervised long-tailed learning. Some methods that attempt to leverage insights from supervised long-tail learning also struggle to form a fitting union with the framework of SSL due to the absence of ground-truth labels [28]. Bridging this paradigmatic gap and efficiently jointing knowledge from SL and SSL represents a promising yet under-explored research direction. Addressing this gap could facilitate deeper synergy between these two learning paradigms and open up new possibilities for developing self-supervised long-tailed learning.
Motivated by the previously mentioned flaws, in this paper, we propose Adaptive Paradigm Synergy (APS). Specifically, we revisit the optimization process of contrastive learning from the perspective of the optimal spatial structure and analyze the effect of different temperature hyper-parameters on the structure of the latent space. Meanwhile, we observe that traditional re-weighting strategy often used in supervised learning can explicitly address their inherent limitations of imbalanced quantity perception. Our strategy of blending these two strategies from different paradigms is what we call dynamic temperature tuning and explicit quantity awareness, driven by a temporary pseudo-label assignment module. Extensive experimental results show that our method effectively integrates these cross-paradigm objectives in a cost-efficient manner with superior performance.
@@
  • We intensively analyze the spatial structural properties of contrastive learning and find that re-balancing strategies from different paradigms (i.e., temperature tuning and re-weighting) can jointly optimize the representations.
  • We present a hybrid cross-paradigm loss called APS which combines dynamic temperature tuning and explicit quantity-aware strategy, bridging the gap between supervised and self-supervised paradigms.
  • We implement comprehensive experiments and give insightful analysis, and the experimental results demonstrate the high superiority and robustness of our approach.
@@ 2 Related Work
Supervised Long-tailed Recognition. Classical supervised long-tail recognition can be categorized in three directions: category-level re-balancing, information augmentation and module improvement [50]. Related to our work, exploration of category-level re-balancing mainly focuses on re-sampling, re-weighting and logit adjustment[25, 55, 26, 44]. Decoupling[25] explores impact of different sampling strategies, and finds that square-root sampling[33] and progressively-balanced sampling are better suited long-tailed datasets. LOCE[14] proposes a memory-agumented feature sampling to dynamically monitor the prediction score of each classes to drive re-sampling rate distribution. Oriented towards supervised contrastive learning paradigm, BCL[55] develops a novel loss for balanced training, which considers the averaging of negative class gradients and the complement of mini-batch tail classes. Logit adjustment[34] utilises training label frequencies to adjust prediction logit. VS loss[26] conducts a comprehensive analysis on impact of additive and multiplicative logit-adjusted losses, and further formulates a novel strategy that combines the best of both worlds.
@@
Explorations on information augmentation seek to introduce additional information into training process, mainly containing transfer learning and data augmentation[46, 51, 17, 29, 35, 15]. SSP[46] systematically analyzes the benefits of combining semi-supervised and self-supervised manners to standard learning for balanced feature space learning. SSD[29] develops a soft labels generator driven by self-supervised distillation, producing a robust auxiliary information to transfer knowledge from long-tailed distribution. DeiT-LT[35] aims to tackle the problems of high efficiency training from scratch, introducing an effective distillation construct from a flat CNN teacher via CLS token and distillation DIST token to support a head expert and a tail expert, respectively. Remix[9] benefits from re-balanced data mixup to boost long-tail learning. SBCL[21] employs subclass-balancing adaptive clustering and bi-granularity contrastive loss to facilitate the implicit mining of sub-classes and optimise the balance of sub-classes.
@@
Module improvement can be divided into classifier design and ensemble learning[11, 51, 23, 45, 49]. Paco[11] introduces a set of parametric class-wise learnable centers to adaptively enhance the optimisation of latent space. RIDE[45] leverages strategy of multiple experts driven by KL-divergence based loss to benefit hard example learning. SADE[49] explores a multi-expert scheme to handle test-agnostic long-tailed recognition, innovating diversity-promoting expertise-guided losses to adaptively aggregate experts for handling unknown distribution.
@@
Self-supervised Long-tailed Recognition expects to automatically correct potential category bias without ground-truth labels and data distributions. There exists several recent explorations for this purpose. SDCLR[22] constructs a pruned dynamic self-competitor for mining forgettable samples, which then are implicitly emphasized in contrastive process. DnC[39] alternates between base contrastive learning and multi-experts hard negative mining, aggregating features from a distillation step to mitigate the collapse of the rare classes. TS[28] leverages a oscillating temperature factor with a cosine schedule to dynamically learn both group-wise and instance-specific features. FASSL[30] presents a frequency-aware prototype learning to identify prototypes guidance reflecting latent distribution from unlabeled data, benefiting re-balanced learning. From a data perspective, BCL-I[54] exploits the memorization effect of DNNs to automatically drive an augmentation of each instance, thereby aiding the learning of discriminative features. GH[53] develops a novel contrasting loss from the perspective of geometric harmonisation, dynamically approaching the class-level uniformity of the latent space.
@@
Pseudo-labels Assignment. DeepCluster[3] leverages the cluster assignments as pseudo-labels, driving an end-to-end unsupervised learning with the soft labels classification pretext task. Subsequently, a large number of methods have been developed to address degenerate solutions, i.e., to prevent collapse. Unlike end-to-end approaches, SCAN[42] adopts a method that decouples representation learning from clustering, ensuring that the clustering process does not rely on low-level features. SeLa[1], on the other hand, focuses on maximizing the information between labels and input data indices. SwAV[5] develops a method that maps images to a set of trainable prototype vectors, which simultaneously avoids the computational cost of pairwise comparisons and improves the model’s scalability and training efficiency.
Figure 2: Model
@@ 3.1 Preliminary
3.2 Temperature Properties for Latent Structure
@@
Latent spatial structure has a direct impact on capability of transferring pre-trained model to other downstream tasks. In recent years, research has been conducted to examine the structure of the latent space for contrastive learning in greater detail [37, 43]. Several evidence from both quantitative and qualitative experiments has demonstrated that temperature parameters exert a profound influence on the uniformity and tolerance of the spatial structure. Previous work [43] calls these two opposing properties as uniformity-tolerance dilemma. To be specific, for Eq. 1 and Eq. 2, given a small $\tau$, the penalty will be dominated by hard negative samples, and therefore more attention will be paid to push the more similar samples farther away, i.e., the tolerance of this feature point is small. Accordingly, when the tolerance of all the points is small, the latent space will be paved over as much as possible, presenting a high uniformity feature. In contrast, a large $\tau$ exhibits high tolerance and low uniformity, therefore feature points will try to occupy as little space as possible.
@@
We give a gradient analysis to illustrate this, for a similarity between $v_{i}$ and a random negative sample $v_{k}$, its gradient is denoted as Eq. 3, and the ratio of its gradient to the sum of gradients with respect to all negative samples is given by Eq. 4. When $sim(\bm{v}_{i},\bm{v}_{k})$ tends to 1, i.e., $v_{k}$ is a hard negative example, with the value of $\tau$ linearly decrease, $r_{k}$ will increase significantly, implying that the hard negative samples gradually dominates learning direction.
$\frac{\partial\mathcal{L}(x_{i})}{\partial s_{i,k}}=\frac{1}{\tau}\frac{S(\bm{v}_{i},\bm{v}_{k},\tau)}{S(\bm{v}_{i},\bm{v}_{i^{\prime}},\tau)+\sum_{v_{j}\in\mathcal{B}_{i}^{-}}S(\bm{v}_{i},\bm{v}_{j},\tau)}$ (3)
$r_{k}=\frac{\exp(sim(\bm{v}_{i},\bm{v}_{k}))/\tau)}{\sum_{v_{j}\in\mathcal{B}_{i}^{-}}\exp(sim(\bm{v}_{i},\bm{v}_{j})/\tau)}$ (4)
3.3 Structure Separability and Category Uniformity
@@
In this section, we give a comprehensive analysis about the expectation of the semantic spacial structure and how dynamic temperature strategy could help reach this ideal structure.
Intuitively, we desire a separable semantic feature to facilitate migration to more downstream tasks [43]. Nevertheless, when we train on a long-tail dataset, due to the potential imbalance of the data, we need a more reasonable structure as target than balanced learning. We describe it as structure separability and category uniformity. As mentioned above, the temperature parameter can control different tolerances for each sample, i.e., in long-tailed learning, we can give different samples different temperatures to form category-level semantic clusters, thereby obtaining a better separable structure. Moreover, due to the inherent imbalance, we also noticed that category uniformity can also help sample to combat tail class neglect and learn robust representation [53]. Specifically, forcing categories with more samples to occupy a similar latent space as categories with fewer samples can effectively combat category representation disparity.
@@
From the perspective of temperature setting, a smaller temperature can make the sample pay more attention to strong negative examples and pursue better separation. On the other hand, stronger uniformity can make the latent space occupied by the category larger, which is more suitable for categories with fewer samples. Conversely, we hope to set a higher temperature for categories with more samples. To this end, we develop a temporary feature clustering module to assign dynamic soft labels as auxiliary information to guide the assignment of temperature. There is no restriction on the clustering method. We use a simple K-means to verify our idea. We will discuss in detail in Sec. 3.5.

### v2

# Dataset Awareness is not Enough: Implementing Sample-level Tail Encouragement in Long-tailed Self-supervised Learning

@@ Abstract
Self-supervised learning (SSL) has shown remarkable data representation capabilities across a wide range of datasets. However, when applied to real-world datasets with long-tailed distributions, performance on multiple downstream tasks degrades significantly. Recently, the community has begun to focus more on self-supervised long-tailed learning. Some works attempt to transfer temperature mechanisms to self-supervised learning or use category-space uniformity constraints to balance the representation of different categories in the embedding space to fight against long-tail distributions. However, most of these approaches focus on the joint optimization of all samples in the dataset or on constraining the category distribution, with little attention given to whether each individual sample is optimally guided during training. To address this issue, we propose Temperature Auxiliary Sample-level Encouragement (TASE). We introduce pseudo-labels into self-supervised long-tailed learning, utilizing pseudo-label information to drive a dynamic temperature and re-weighting strategy. Specifically, We assign an optimal temperature parameter to each sample. Additionally, we analyze the lack of quantity awareness in the temperature parameter and use re-weighting to compensate for this deficiency, thereby achieving optimal training patterns at the sample level. Comprehensive experimental results on six benchmarks across three datasets demonstrate that our method achieves outstanding performance in improving long-tail recognition, while also exhibiting high robustness.
@@ 1 Introduction
Figure 1: Comparison between SimCLR [5] baseline trained on balanced CIFAR10[25] and imbalanced CIFAR10-LT [11] and our TASE trained on CIFAR10-LT. Three Splits (Head, Mid and Tail) are shown as different colours (blue, red and green) respectively [29]. We also report the average accuracy and range of variation over ten classes.
@@
Self-supervised learning (SSL) [17, 18], which aims to learn robust and generic feature representations from unlabeled data, has been consistently explored in recent years [7, 3, 45]. Contrastive learning, a leading SSL paradigm, seeks to learn discriminative representations by encouraging the embedding similarity of augmented view pairs, and has achieved remarkable success across several domains [46, 15]. Nevertheless, most contrastive learning methods are conducted on curated datasets like ImageNet [12]. In contrast, real-world data distributions are mainly characterized by long-tailed distributions [34], which could cause a significant degradation in performance, as illustrated in Fig. 1. On the relatively simple CIFAR10-LT dataset, the baseline model demonstrates a notable performance collapse. For certain tail classes, the accuracy even falls below 30%, indicating that SSL still suffers from the long-tailed problem. The difficulty of detecting or modeling such implicit imbalance is a significant obstacle to migrating SSL to large-scale, uncurated, real-world datasets.
In recent years, for the supervised paradigm, the long-tailed problem has been extensively studied from multiple perspectives, deriving strategies such as re-weighting [53], re-sampling, and ensemble learning, which have achieved significant success. In terms of re-weighting, a common approach is to increase the weight proportion of tail-class samples, in order to occupy a larger gradient and encourage the model to focus on tail-class samples. Turning to self-supervised learning, particularly contrastive learning, some works have proposed using temperature assignment to counteract the effects of long-tailed distributions [41, 26]. Specifically, models can be made to focus on different semantics by adjusting the temperature parameters of the contrastive objective. A lower temperature encourages sample features to push similar samples away as much as possible, independently forming semantic clusters. This is a pattern emphasizes category discrimination, while a higher temperature allows greater tolerance for similar samples, focusing more on semantic difference learning and instance discrimination. Such learning pattern difference is a potential way to combat long-tailed distributions.
Then our perspective shifts back to self-supervised long-tailed learning, some research has sought to adapt the aforementioned mechanisms to this domain. TS [26] attempts to employ a global temperature with cosine variation to modify the model’s learning pattern between instance discrimination and category discrimination. GH [52] emphasizes the necessity for spatial homogeneity in imbalanced learning scenarios and proposes a plug-and-play objective for harmonizing the embedding space of each category. Nevertheless, none of these efforts achieve optimal training for all samples, and intuitively, it is more efficient to ensure that each sample is appropriately trained than to constrain the global.
Motivated by the above considerations, we notice pseudo-labels assignment in semi-supervised learning, and then use pseudo-labels of each sample to achieve sample-level encouragement. Nevertheless, due to the long-tailed distribution, the tail classes with few samples are more likely to yield severe misassignments. In order to avoid serious misleading of the tail classes, several pseudo-labels assignment methods from semi-supervised learning, which use pseudo-labels as targets for consistency regularization [36, 47], cannot be directly transferred to the long-tailed scenario. Therefore, we make the following attempts and propose TASE. Specifically, we assign appropriate temperature based on the category to which the pseudo-labels of samples belong. This allows the head class to learn semantic differences from other similar samples, and while the tail class is separated from other sample clusters, thereby generating individual categories. Nevertheless, during the training process, the head and tail classes are assigned the maximum and minimum temperatures respectively, but the range of temperature parameters is often treated as a fixed parameter, which results in the strength of tail encouragement having the same intensity for datasets with different imbalance ratios. We call this as quantity awareness flaw of temperature strategy. To address this flaw, we propose re-weighting strategy to provide explicit quantity awareness. In particular, we add a weight to each sample that is inversely proportional to the number of samples in the category to which it belongs.
Compared to existing work, our method has the following advantages: 1) Compared to previous work on TS [26] and GH [52], we are able to modify the training process at the sample-level, providing the most appropriate encouragement to each sample to counteract the long-tailed distribution. 2) Compared to the approach in semi-supervised learning, where pseudo-labels are directly used as targets for consistency regularization, our method does not require pseudo-labels for direct training, and it is better at resisting the misleading effects of incorrect label assignments.
@@
  • We note the drawback of existing work on long-tail self-supervised learning that it is difficult to perform tail encouragement for the sample level, and propose to introduce the idea of pseudo assignment into long-tail self-supervised learning to perform sample-specific learning for each sample.
  • We propose TASE as a preliminary attempt to implement sample-level control in long-tail self-supervised learning, assigning appropriate temperatures driven by pseudo assignments for different samples during training to achieve optimal training patterns.
  • We implement comprehensive experiment across six benchmarks on three commonly used datasets, and the experimental results demonstrate our model achieves superior performance.
@@ 2 Related Work
Supervised Long-tailed Recognition. Classical supervised long-tail recognition can be categorized into three directions: category-level re-balancing, information augmentation and module improvement [49]. Related to our work, exploration of category-level re-balancing mainly focuses on re-sampling, re-weighting and logit adjustment [21, 53, 24, 42]. Decoupling [21] explores the impact of different sampling strategies, and finds that square-root sampling [31] and progressively-balanced sampling are better suited for long-tailed datasets. LOCE [13] proposes a memory-agumented feature sampling to dynamically monitor the prediction score of each classes to drive re-sampling rate distribution. Oriented towards the supervised contrastive learning paradigm, BCL [53] develops a novel loss for balanced training, which considers the averaging of negative class gradients and the complement of mini-batch tail classes. Logit adjustment [32] utilises training label frequencies to adjust the prediction logit. VS loss [24] conducts a comprehensive analysis on the impact of additive and multiplicative logit-adjusted losses, and further formulates a novel strategy that combines the best of both worlds.
@@
Explorations on information augmentation seek to introduce additional information into training process, mainly containing transfer learning and data augmentation [44, 50, 16, 27, 33, 14]. SSP [44] systematically analyzes the benefits of combining semi-supervised and self-supervised manners to standard learning for balanced feature space learning. SSD [27] develops a soft labels generator driven by self-supervised distillation, producing robust auxiliary information to transfer knowledge from long-tailed distribution. DeiT-LT [33] aims to tackle the problems of high efficiency training from scratch, introducing an effective distillation construct from a flat CNN teacher via CLS token and distillation DIST token to support a head expert and a tail expert, respectively. Remix [8] benefits from re-balanced data mixup to boost long-tail learning. SBCL [19] employs subclass-balancing adaptive clustering and bi-granularity contrastive loss to facilitate the implicit mining of sub-classes and optimise the balance of sub-classes.
@@
Module improvement can be divided into classifier design and ensemble learning [10, 50, 22, 43, 48]. Paco [10] introduces a set of parametric class-wise learnable centers to adaptively enhance the optimisation of latent space. RIDE [43] leverages a strategy of multiple experts driven by KL-divergence based loss to benefit hard example learning. SADE [48] explores a multi-expert scheme to handle test-agnostic long-tailed recognition, innovating diversity-promoting expertise-guided losses to adaptively aggregate experts for handling unknown distribution.
@@
Self-supervised Long-tailed Recognition expects to automatically correct potential category bias without ground-truth labels and data distributions. There are several recent explorations for this purpose. SDCLR [20] constructs a pruned dynamic self-competitor for mining forgettable samples, which are then implicitly emphasized in the contrastive process. DnC [38] alternates between base contrastive learning and multi-experts hard negative mining, aggregating features from a distillation step to mitigate the collapse of the rare classes. TS [26] leverages an oscillating temperature factor with a cosine schedule to dynamically learn both group-wise and instance-specific features. FASSL [28] presents a frequency-aware prototype learning to identify prototypes guidance reflecting latent distribution from unlabeled data, benefiting re-balanced learning. From a data perspective, BCL-I [51] exploits the memorization effect of DNNs to automatically drive an augmentation of each instance, thereby aiding the learning of discriminative features. GH [52] develops a novel contrasting loss from the perspective of geometric harmonisation, dynamically approaching the class-level uniformity of the latent space.
@@
Pseudo-labels Assignment. DeepCluster [2] leverages the cluster assignments as pseudo-labels, driving an end-to-end unsupervised learning with the soft labels classification pretext task. Subsequently, a large number of methods have been developed to address degenerate solutions, i.e., to prevent collapse. Unlike end-to-end approaches, SCAN [40] employs a method that decouples representation learning from clustering, ensuring that the clustering process does not depend on low-level features. SeLa [1], on the other hand, focuses on maximizing the information between labels and input data indices. SwAV [4] develops a method that maps images to a set of trainable prototype vectors, which simultaneously avoids the computational cost of pairwise comparisons and improves the model’s scalability and training efficiency.
@@ 3.1 Preliminary
3.2 Temperature Property
@@
In recent years, research has been conducted to examine the mechanisms of different temperature parameters in greater details [35, 41]. Several evidence from both quantitative and qualitative experiments has demonstrated that temperature parameters exert a profound influence on the uniformity and tolerance of training process. Previous work [41] calls these two opposing properties as the uniformity-tolerance dilemma, corresponding to instance-discrimination and category-discrimination, respectively. To be specific, for Eq. 1 and Eq. 2, given a small $\tau$, the penalty will be dominated by hard negative samples, and therefore more attention will be paid to push the more similar samples farther away, i.e., the tolerance of this feature point is small. Accordingly, when the tolerance of all the points is small, the latent space will be paved over as much as possible, presenting a high uniformity feature. In contrast, a large $\tau$ exhibits high tolerance and low uniformity; therefore, feature points will try to occupy as little space as possible.
@@
We give a gradient analysis to illustrate this. The gradient for the similarity between $v_{i}$ and a random negative sample $v_{k}$ is denoted as Eq. 3, and the ratio of its gradient to the sum of gradients with respect to all negative samples is given by Eq. 4. When $sim(\bm{v}_{i},\bm{v}_{k})$ tends to 1, i.e., $v_{k}$ is a hard negative example, with the value of $\tau$ linearly decrease, $r_{k}$ will increase significantly, implying that the hard negative samples gradually dominates gradient direction.
$\frac{\partial\mathcal{L}(x_{i})}{\partial sim(\bm{v}_{i},\bm{v}_{k})}=\frac{1}{\tau}\frac{S(\bm{v}_{i},\bm{v}_{k},\tau)}{S(\bm{v}_{i},\bm{v}_{i^{\prime}},\tau)+\sum_{\bm{v}_{j}\in\mathcal{B}_{i}^{-}}S(\bm{v}_{i},\bm{v}_{j},\tau)}$ (3)
$r_{k}=\frac{\exp(sim(\bm{v}_{i},\bm{v}_{k}))/\tau)}{\sum_{\bm{v}_{j}\in\mathcal{B}_{i}^{-}}\exp(sim(\bm{v}_{i},\bm{v}_{j})/\tau)}$ (4)
3.3 Structural Separability and Category Uniformity
@@
In this section, we provide an analysis about the expectation of the semantic spatial structure and how the dynamic temperature strategy could help reach this ideal structure.
In order to obtain robust downstream task performance, higher requirements are demanded for the semantic space structure of the data representation, which we describe as structural separability and category uniformity. As shown in Fig. 2, structural separability expects that samples from different categories form category semantic clusters and maintain strong linear separability. At the same time, in order to focus on the tail categories, category uniformity expects that each category occupying embedding space should be similar, irrespective of the number.
@@
From the perspective of temperature setting, a lower temperature can make the sample pay more attention to strong negative examples and pursue better separation. On the other hand, stronger uniformity can make the latent space occupied by each category larger, which is more suitable for categories with fewer samples. Conversely, we hope to set a higher temperature for categories with more samples, encouraging the model to learn the inherent semantic differences between samples and maintain high generalization across a variety of transfer learning tasks. To this end, we utilize a temporary feature clustering module to assign dynamic pseudo-labels as auxiliary information to guide the assignment of temperature. If the sample belongs to the head class, a higher temperature is set, and if it belongs to the tail class, a lower temperature is assigned. There is no restriction on the clustering method. We use a simple K-means to verify our idea.
Figure 2: Illustration of the structure of the desired embedding space. We use different icons in different colors to represent the samples from different categories and indicate the size of the space occupied by each category. We expect the clusters to be at a certain distance from each other and to occupy a similar amount of space.